OpenStudy (anonymous):

What is 3 cubes?? 4 or 27

OpenStudy (diyadiya):

27

OpenStudy (anonymous):

3 cube = 3 * 3 * 3 = 27

OpenStudy (diyadiya):

3 x 3 x 3 =27 =3^3

OpenStudy (anonymous):

What he/she said ^ :)

OpenStudy (liliakarina):

3 * 3 * 3= 27